OpenStudy (anonymous):

What is the y-intercept in this linear equation? y=-4x+4

OpenStudy (anonymous):

4

OpenStudy (anonymous):

What is the x-intercept in this linear equation? 2y-4x=-12

OpenStudy (anonymous):

x-intercept = 3 and y-intercept = -6

OpenStudy (anonymous):

What is the y-intercept for this linear equation? 6x-y=9

OpenStudy (anonymous):

The equation here given is y = -4x + 4 Lets solve and make the equation in the form of ax +by +c = 0 Thus, the equation becomes 4x + y = 4 Now, we can compare the equation with ax + by + c = 0 Then a = 4 b = 1 c = 4 Now, we can use formula to find out the x-intercept x-intercept = c/a = 4/4 = 1 Thus, the x-intercept is 1